Bell Company, a manufacturer of audio systems, started its production in October 2020. For the preceding 3 years, Bell had been a retailer of audio systems. After a thorough survey of audio system markets, Bell decided to turn its retail store into an audio equipment factory.

Raw material costs for an audio system will total $77 per unit. Workers on the production lines are on average paid $14 per hour. An audio system usually takes 5 hours to complete. In addition, the rent on the equipment used to assemble audio systems amounts to $6,000 per month. Indirect materials cost $5 per system. A supervisor was hired to oversee production; her monthly salary is $4,000.

Factory janitorial costs are $1,800 monthly. Advertising costs for the audio system will be $9,200 per month. The factory building depreciation expense is $6,000 per year. Property taxes on the factory building will be $9,600 per year.

Part 1

Assuming that Bell manufactures, on average, 1,000 audio systems per month, enter each cost item on your answer sheet, placing the dollar amount per month under the appropriate headings. Total the dollar amounts in each of the columns.

Product Costs


Cost Item

Direct
Materials

Direct
Labor

Manufacturing
Overhead

Period
Costs

Raw materials

$

$

$

$

Wages for workers
Rent on equipment
Indirect materials
Factory supervisor’s salary
Janitorial costs
Advertising
Depreciation on factory building
Property taxes on factory building

$

$

$

$

eTextbook and Media

Save for Later

Attempts: 0 of 3 used

Submit Answer

Using multiple attempts will impact your score.

10% score reduction after attempt 2

Part 2

Compute the cost to produce one audio system. (Round answer to 2 decimal places, e.g. 15.25.)

Production cost per system
Answer #1
Product Costs

Cost Item
Direct
Materials
Direct
Labor
Manufacturing
Overhead
Period
Costs
Raw materials $77000 $ $ $
Wages for workers 70000
Rent on equipment 6000
Indirect materials 5000
Factory supervisor’s salary 4000
Janitorial costs 1800
Advertising 9200
Depreciation on factory building 500
Property taxes on factory building 800
$77000 $70000 $18100 $9200

Total cost to produce audio systems

= Direct material + direct labor + manufacturing overhead

It won't include advertising costs

So cost = 77000+70000+18100 = $165100

Cost per audio system = 165100/1000 = $165.10